# Advancing African-Accented English Speech Recognition: Epistemic Uncertainty-Driven Data Selection for Generalizable ASR Models

### v6

@@ 5 Conclusion
We combined several AL paradigms, the CSA, and EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced the U-WER metric to track model adaptation to hard accents. Our experiments showed a remarkable 27% WER ratio improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This demonstrates our approach’s efficiency and potential to lower the barriers to ASR technologies in underserved regions significantly. Our method enhances model robustness and generalization across various domains, datasets, and accents, which are crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
@@ 6 Limitations
In discussing trade-offs (Section 4), we noted that while our approach enhances performance, particularly with linguistically rich accents, a stopping criterion is essential for complex domains like the clinical one, to balance adaptation rounds with the pool size. With better resources, we would consider implementing Deep Ensembles [30] as an alternative to our current MC-Dropout method for estimating epistemic uncertainty, and leverage other acquisition functions (such as BALD, BatchBALD) highlighted in this work.
Acknowledgments and Disclosure of Funding
The author is very grateful to Ines Arous for her help in reviewing the manuscript and suggesting very important changes to improve the quality of the paper. The author acknowledges the support of the Mila Quebec AI Institute for computing resources.
The author acknowledges Intron Health for providing the Afrispeech-200 dataset. The author is grateful to Atnafu Lambebo Tonja, Chris Chinenye Emezue, Tobi Olatunji, Naome A Etori, Salomey Osei, Tosin Adewumi, and Sahib Singh; for their help in the early stage of this project.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] (See Section 4)
    2. (b)
      Did you describe the limitations of your work? [Yes] (See Section 6)
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] (See the abstract)
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] (See Tables 2, 1, A.1)
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] (See Section 4, and Figures 2, 3 and 4)
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] (See Section 3.3)
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] (See Section 3.1)
    2. (b)
      Did you mention the license of the assets? [Yes] (See Section 3.1)
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [N/A]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
@@ A.2 Country Statistics
Table 6 shows the statistics of the countries across the AfriSpeech-200 dataset.

### v7

@@ 5 Conclusion
We combined several AL paradigms, the CSA, and the EU to create a novel multi-round adaptation process for high-performing pretrained speech models, aiming to build efficient African-accented English ASR models. We introduced the U-WER metric to track model adaptation to intricate accents. Our experiments demonstrated a remarkable 27% WER ratio improvement while reducing the data required for effective training by approximately 45% compared to existing baselines. This reflects the efficiency and potential of our approach to lower the barriers to ASR technologies in underserved regions significantly. Our method enhances model robustness and generalization across various domains, datasets, and accents, which are crucial for scalable ASR systems. This also helps mitigate bias in ASR technologies, promoting more inclusive and fair AI applications.
@@ 6 Limitations
In discussing trade-offs (Section 4), we noted that while our approach enhances performance, particularly with linguistically rich accents, a stopping criterion is essential for complex domains like the clinical one to balance adaptation rounds with the pool size. With better resources, we would consider implementing Deep Ensembles (Lakshminarayanan et al. (2017)) as an alternative to our current MC-Dropout method for estimating epistemic uncertainty and leveraging other acquisition functions (such as BALD, BatchBALD) highlighted in this work.
7 Acknowledgments
The authors are very grateful to Prof. Ines Arous for her help in reviewing the manuscript and suggesting very important changes to improve the quality of the paper. The author acknowledges the support of the Mila Quebec AI Institute for computing resources.
The authors acknowledge Intron Health for providing the Afrispeech-200 dataset. The authors are grateful to Atnafu Lambebo Tonja, Chris Chinenye Emezue, Tobi Olatunji, Naome A Etori, Salomey Osei, Tosin Adewumi, and Sahib Singh for their help in the early stage of this project.
@@ A.2 Country Statistics
Table 6 shows the countries’ statistics across the AfriSpeech-200 dataset.